OpenStudy (anonymous):

A woman with her hands outstretched has a moment of inertia of 2 kg m2. She rotates on a stool at an angular velocity of 6 rad s-1. At a certain instant she grabs a 3 kg object in each hand. Holding them 0.8 m from the axis of rotation. If the objects are initially at rest, what is her angular velocity??

OpenStudy (anonymous):

Are you familiar with Steiner's theorem?

OpenStudy (anonymous):

I get: I' = mr^2 = (3kg+3kg)* (0,8)^2 = 3,84 kgm2 w' = (Iw) / I' = (2*6)/3,84 = 3.125 rad/sek But the right answrer is 2.02 rad/sek

OpenStudy (anonymous):

No